Suppose that on the average, 7 students enrolled in a small liberal arts college have their automobiles stolen during the semester. What is the probability that less than 3 students will have their automobiles stolen during the current semester? Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

mean = = 7

Using poisson probability formula,

P(X = x) = (e- * x ) / x!

P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2)

= (e-7 * 70) / 0! + (e-7 * 71) / 1! + (e-7 * 72) / 2!

= 0.02964

The probability that less than 3 students will have their automobiles stolen during the current semester is = 0.0296